OpenStudy (anonymous):

A loaded train travelling at 100 m/s can decelerate at 20 m/s2. How far in advance must the brakes be applied to come to a stop?

OpenStudy (anonymous):

Initial Velocity of the train(u)=100 m/s Deceleration(a)=-(a) =-20m/s^2 And Final velocity(v)=0m/s We know that, Time(t)= v-u/t =(0-100)/-20 =-100/-20 =5 sec. Now, v^2=u^2+2as where 's' denotes distance till where the train travels, or, 0=(100)^2+ 2*(-20)*s or, 0= 10000-40s or, 40s=10000 or, s= 10000/40 s= 250 m. Hence, the brakes must be applied at 250 metre in advance.

OpenStudy (anonymous):

sorry...its t=v-u/a